# Kilohertz Gravitational Waves from Binary Neutron Star Mergers: Full Spectrum Analyses and High-density Constraints on Neutron Star Matter

### v1

@@ B.1 Bayesian framework
When $\mathcal{B}_{\rm{A}}^{\rm{B}}>1$ ($<1$), the model hypothesis B (A) is favored. Furthermore, we introduce the SNR from Eq. (18) as
@@ B.2 Relative binning
Since this ratio is smooth, it can be approximated by a piecewise linear function Zackay et al. (2018):
@@
where $b$ represents a frequency bin and $f_{m}$ is the mean frequency of the bin. Some components of the inner products involved in the likelihood computation can be precomputed to improve the efficiency:
@@
Relative binning technique is effective for waveform analyses up to the merger part, as the frequency domain representations are smooth in this regime. Instead, in the PM regime, the waveform structure becomes highly complex and rapidly varying in frequency, rendering relative binning inefficient and insufficiently accurate. Nevertheless, NRPMw is fast, being phenomenological, therefore acceleration techniques are not critical as in the case of TEOBResumSPA. In our full spectrum analyses, we thus employ relative binning only in the IM portion and we perform standard PE for the PM part of the spectrum. To implement this approach, we introduce cutoff frequency $f_{\rm cut}=2048~{}{\rm Hz}$. The likelihood is computed separately for the two regimes: using relative binning for frequencies below $f_{\rm cut}$ (IM) and using the full, unaccelerated Bayesian PE for frequencies above the cutoff (PM). The total likelihood is obtained by summing the two contributions, cfr. Eq (8). This hybrid approach enables efficient and accurate PE for GW signals spanning both the IM and PM regimes.
The pre-computation of some terms and the evaluation of the waveforms only on the bin edges enable PE of coalescing compact binaries in a very short amount of time: it is possible to run a PE on a single CPU in 44 minutes for a binary black holes, using the TEOBResumS template, and in 275 minutes for a BNS with TaylorF2. In the case of full spectrum BNSs waveforms with TEOBResumSPA_NRPMw, we run on a cluster with 192 CPUs in about 40 hours, making use of the MPI-parallelization of bajes Breschi et al. (2021).
@@ Appendix C Results
We report in Table LABEL:tab:IM2 the results of the injection-recovery with IM-only TEOBResumSPA template relative to Fig. 4 and 5.

### v2

@@ B.1 Bayesian framework
When $\mathcal{B}_{\rm{A}}^{\rm{B}}>1$ (<1<1), the model hypothesis B (A) is favored. Furthermore, we introduce the SNR from Eq. (18) as
@@ B.2 Relative binning
Since this ratio is smooth, it can be approximated by a piecewise linear function Zackay et al. (2018)
@@
where $b$ represents a frequency bin and $f_{m}$ is the mean frequency of the bin. Some components of the inner products involved in the likelihood computation can be precomputed to improve the efficiency
@@
Relative binning technique is effective for waveform analyses up to the merger part, as the frequency domain representations are smooth in this regime. Instead, in the PM regime, the waveform structure becomes highly complex and rapidly varying in frequency, rendering relative binning inefficient and insufficiently accurate. Nevertheless, NRPMw is fast, being phenomenological, therefore acceleration techniques are not critical as in the case of TEOBResumSPA. In our full spectrum analyses, we thus employ relative binning only in the IM portion and we perform standard PE for the PM part of the spectrum. To implement this approach, we introduce cutoff frequency $f_{\rm cut}=2048~{\rm Hz}$. The likelihood is computed separately for the two regimes: using relative binning for frequencies below $f_{\rm cut}$ (IM) and using the full, unaccelerated Bayesian PE for frequencies above the cutoff (PM). The total likelihood is obtained by summing the two contributions, cf. Eq (8). This hybrid approach enables efficient and accurate PE for GW signals spanning both the IM and PM regimes.
The precomputation of some terms and the evaluation of the waveforms only on the bin edges enable PE of coalescing compact binaries in a very short amount of time: it is possible to run a PE on a single CPU in 44 minutes for a binary black holes, using the TEOBResumS template, and in 275 minutes for a BNS with TaylorF2. In the case of full spectrum BNSs waveforms with TEOBResumSPA_NRPMw, we run on a cluster with 192 CPUs in about 40 hours, making use of the MPI-parallelization of bajes Breschi et al. (2021).
@@ Appendix C Results
We report in Table LABEL:tab:IM2 the results of the injection recovery with IM-only TEOBResumSPA template relative to Fig. 4 and 5.